OpenStudy (anonymous):

Could somebody explain to me why the following statement is true, " arctan(1) = pi/4"

terenzreignz (terenzreignz):

Do you know what arctan is?

Parth (parthkohli):

What is pi/4? It is 90 degrees. Now check if \(\tan 90^{\circ} = 1\)

OpenStudy (anonymous):

Yes, I understand that it is the inverse of the tangent function.

Parth (parthkohli):

Is \(\tan90^{\circ} = 1\)?

terenzreignz (terenzreignz):

ParthKohli can put it much more simply than I can T.T

Parth (parthkohli):

haha

OpenStudy (anonymous):

Isn't tan at 90 undefined?

Parth (parthkohli):

In other words, \(\tan = \large {\sin \over \cos}\)

Parth (parthkohli):

Yes it is undefined then ^_^

terenzreignz (terenzreignz):

pi/4 is 45 degrees o.o

OpenStudy (anonymous):

** tan 45, he means.

Parth (parthkohli):

OMG SORRY

Parth (parthkohli):

\( \color{Black}{\Rightarrow \tan45^{\circ} = 1}\)

terenzreignz (terenzreignz):

This is awkward...

OpenStudy (unklerhaukus):



Parth (parthkohli):

I guess you already know the important ratios, do you?

terenzreignz (terenzreignz):

Don't you guys think we're complicating things a bit?

Parth (parthkohli):

Haha no. \(\sin 90 = 1\) \( \color{Black}{\Rightarrow\cos 0 = 1 }\) \( \color{Black}{\Rightarrow \tan 45 = 1 }\)

OpenStudy (unklerhaukus):

\[\arctan\left(\frac{\text {opposite}}{\text{adjacent}}\right) = \theta\] \[\arctan\left(\frac 11\right) =45°= \pi/4\]

OpenStudy (anonymous):

The biggest issue is that I am completely knew to trig and like when I learned algebra, I needed steps to find my solutions. Could somebody give me some step-by-step instructions to solving the problem?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \tan 45 = \large {\sin 45 \over \cos 45 } = {\Large {1 \over \sqrt2} \over {1 \over \sqrt2}} }\)

terenzreignz (terenzreignz):

Well, you know the relationship between functions and their inverses?

Parth (parthkohli):

That leads us to the statement tan 45 = 1

OpenStudy (anonymous):

Yes, I understand that (for lack of a better explanation) the domain and range is flipped.

Parth (parthkohli):

Well you can refer to @UnkleRhaukus 's diagram

Parth (parthkohli):

(90,1) becomes (1,90) where 90 = x and f(x) is sin.

Parth (parthkohli):

Similarly, (45,1) becomes (1,45) that's it ^_^

terenzreignz (terenzreignz):

other than that, for instance, Let f be a bijective* function, such that g = f^-1 then if a is in the range of f, then what's f(g(a)) ?

OpenStudy (unklerhaukus):

if your arc tan is one it means the opposite and adjacent sides are equal. the sum of the angles in a triangle is 180° the triangle has a right angle , because we are using trig ratios , and the angles are equal because the opp and adjacent are equal 180°-90°=2x45°

OpenStudy (anonymous):

\[\tan45 = 1\] Or,\[\tan(\pi /4) = 1\] Because, \[45 = \pi /4\] So, \[\tan^{-1} (1) = \tan^{-1} (\tan \pi /4) = \pi /4\] is the answer...

terenzreignz (terenzreignz):

Long story, short, if you compose a function with its inverse, they'd cancel each other out :) so since tan (pi/4) = 1 then arctan[tan (pi/4)] = arctan 1 since arctan is the inverse of tan, they'd cancel out so... pi/4 = arctan 1